Tyrion:

The aquarium has 3 more yellow fish than green fish. 60 percent of the fish are yellow. How many green fish are in the aquarium? Show your work.

snowflake0531:

choose the variables first

snowflake0531:

Yellow fish and green fish, x and y

snowflake0531:

aka, yellow fish is x, green fish is y

snowflake0531:

How would you setup the equation for 3 more yellow fish than green fish?

snowflake0531:

Should i take that as you don't know how?

snowflake0531:

Well, since there are 3 more yellow fish, green plus 3, equals yellow so, y=x+3 agreed?

Tyrion:

ok so y=x+3 bekause there are 3 more yellow fish than green,,got it karry on

snowflake0531:

. 60 percent of the fish are yellow. 0.6(x+y) = x Yea?

snowflake0531:

I'm gonna fish this then ig lol

snowflake0531:

and guess what, I'm changing the variables, the confuse me Yellow will by y, green will be g

snowflake0531:

The aquarium has 3 more yellow fish than green fish. 60 percent of the fish are yellow. How many green fish are in the aquarium? Show your work.

snowflake0531:

So, g+3=y and (g+y)0.6 = y

snowflake0531:

The easiest way would be desmos, which would be https://www.desmos.com/calculator/n5ezfcz4wh

snowflake0531:

And because you want the green, (desmos made me change it to x), it is 6, 6 green fish

snowflake0531:

Or, we could try the substitution method

snowflake0531:

g+3=y and (g+y)0.6 = y g+3 = (g+y)0.6 g+3 = (g+g+3)0.6 g+3 = 1.2g + 1.8 3 = 0.2 g + 1.8 1.2 = 0.2g g=6 6 fish

snowflake0531:

So, there are 6 fish, you can use either way, although I wouldn't recommend the elimination method for this problem